From nch@roe.ac.uk Fri Dec 3 18:20:16 2004 Date: Fri, 3 Dec 2004 12:52:52 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Harvey MacGillivray , Ian Bond , Mike Read , Nigel Hambly , Bob Mann Cc: CCs for WSA weekly meeting minutes distribution -- Clive Davenhall , Andrew Lawrence , Andy Adamson , Peter Shillan , John Taylor , Jim Emerson , Malcolm Stewart , Martin Hill , Mike Irwin , Peredur Williams Subject: WFAU WSA weekly meeting minutes, 3rd December 2004 Minutes of WFCAM Science Archive meeting: 3rd December 2004 ------------------------------------------------------------- ------------------------------------------------------------- Present: NCH, PMW, JMS, ETWS, MAR Apologies: JPE, RGM, MCH, AL, HMG, JDT, GPS, DONM: 10am, Friday 10th December, plate library Actions discharged: ------------------- ACTION: PMW to clarify what remains of development hardware money with JPE. Discharged: it seems there is probably of order 30K still available. CORRECTION: Apparently there is not 30K of VDFS money available for hardware. ACTION: ETWS to liaise with MJI and JRL to test out the CASU/WFAU VDFS interface. Discharged; awaiting JRL's return from Hawaii with a tape of engineering frames. ACTION: NCH to contact Jan at JHU to enquire as to the possibility of getting DR3 catalogues/spectra directly (as we did for DR2 cats). Discharged; JHU have been contacted but no response as yet (possibly due Thanksgiving holidays last week) ACTION: AL to email AA to enquire as to status of deliberations on user registration for the WSA. Discharged; AA has been contacted; awaiting reply. Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR to check out SQL Server networked server functionality for cross-server querying. Continues; NCH suggested we should do some experiments with the linked SQL Servers ahmose & amenhotep once the dust has settled on their reorganisation over the next couple of weeks. ACTION: NCH to remember to action a team programmer to implement range checking in the low-level ingest codes. ETWS & NCH have been discussing this; likely that the new developer will be assigned this task to start them off. ACTION: JDT to liaise with MCH concerning an end-to-end test of the PAL deployment. Continues; see Astrogrid Deployment below. ACTION: PMW to apply a little pressure to the premises team to solve the continuing problem of heat overload in C1. Continues; a quote has been received by Premises for an extra A/C unit. No word as yet on when the purchase/installation may happen... ACTION: ALL to review risk register and think of any new internal or external risks that should be documented in the risk register. Defered until top level plan is explicitly defined. PMW tabled the VISTA risk register as an example. ACTION: JPE to design and set up centralised VDFS web pages. Progressing. (there is now a link to it from our WSA Twiki). - continuing; thanks for progressing these. Actions carried forward from 29/10/04 meeting: ---------------------------------------------- None this week. Specific points and new actions: -------------------------------- Project management: The team noted the upcoming VDMT meeting on the 6th Dec. PMW will assemble the usual table stats on progress; NCH will assemble some associated narrative. JMS requested a meeting with NCH & PMW to thrash out the top-level project plan for Vista science archiving work. ACTION: NCH, JMS & PMW to meet Wednesday 8th Dec at 10am in the Plate Library to thrash out the VDFS top level forward-plan of work. Comments and issues arising from CASU fortnightly minutes: No new minutes as of 3/12/04 am. Networking: Nothing new to report. Hardware: Good news: the reconfigured server "amenhotep" is now live online with an enhanced SSA that now includes SDSS DR2. The old server "ahmose" will be kept online until after the Astrogrid demos (week after next) just in case of any problems with either server. NCH will arrange for Eclipse to come in after that to reconfigure ahmose along the same lines as amenhotep to create the new offline load server. ACTION: NCH to contact Eclipse to arrange reconfiguration of old catalogue server ahmose. Software: NCH and ETWS have been working on CU19/20: some small changes to curation metadata tables in the WSA and refactoring of code. ETWS has also enhanced parsing scripts for both schema and DB table constraints, and has been working on integrating CASU pixel processing tools (ie. CU13/14). Otherwise, MAR & ETWS have been attending the NeSC datamining course. SSA: No news other than redeployment of the archive on a server with a new disk subsystem, and incorporation of SDSS DR2. Astrogrid deployment: Important now to maintain a stable full SSA for the Astrogrid demos the week after next, hence both catalogue servers will hold a full SSA and should be tested under PAL access. Miscellaneous: CORRECTION: it was JDT that brought the biccies to the last meeting, not JNTD.